OpenStudy (anonymous):

Steve has the following scores on the exams for the semester: 96, 85, and 89. Determine the minimum score Steve needs to make on the fourth exam to have a final average of an “A” (a 93 or better).

OpenStudy (anonymous):

(96 + 85 + 89 + x )/4 = 93 (270 + x)/4 = 93 (now multiply the equation by 4 270 + x = 372 x = 372 - 270 x = 102 check... (96 + 85 + 89 + 102)/4 = 93 372/4 = 93 93 = 93 (correct) I got 102, but you might need to get a second opinion

OpenStudy (anonymous):

ok thanks